OpenStudy (anonymous):

Can someone please help me figure out how to set up this problem? Consider the points P such that the distance from P to A(-1, 5, 3) is twice the distance from P to B(6, 2, -2). Show that the set of all such points is a sphere, and find its center and radius.

OpenStudy (dumbcow):

Use the distance formula Let point P be (x,y,z) \[\sqrt{(x+1)^{2}+(y-5)^{2}+(z-3)^{2}} = 2*\sqrt{(x-6)^{2}+(y-2)^{2}+(z+2)^{2}}\] square both sides \[(x+1)^{2}+(y-5)^{2}+(z-3)^{2} = 4*((x-6)^{2}+(y-2)^{2}+(z+2)^{2})\] Combine like terms on one side, put constants on other side Use completing the square to get it in standard form \[(x-\frac{25}{3})^{2}+(y-1)^{2}+(z+\frac{11}{3})^{2} = \frac{1178}{9}\]

OpenStudy (dumbcow):

correction, right side should be (332/9)

OpenStudy (anonymous):

thanks so much!

OpenStudy (dumbcow):

your welcome